OpenStudy (anonymous):

Simplify y^-10. Answer (a) −y^10 (b) 1/Y^10 (c) 1/y^-10 (d) -y

OpenStudy (anonymous):

by power of exponent : it will be b option

OpenStudy (anonymous):

i thnk is a

OpenStudy (campbell_st):

\[y^{-10} = \frac{1}{y^{10}}\] hope this helps

OpenStudy (campbell_st):

negative indices are for fractions

OpenStudy (lgbasallote):

\[\huge a^{-m} = \frac{1}{a^m}\]

OpenStudy (anonymous):

the question is saying y to the power -10 which is 1/y^10

OpenStudy (anonymous):

\[y^{-10}=1/y ^{10}\]

OpenStudy (anonymous):

put that in log form hehehe

OpenStudy (anonymous):

help her. just say the answer is B